An investment doubles over a period of six years. Determine the simple interest rate


1
Expert's answer
2021-10-26T08:13:47-0400

Simple interest


"A=P(1+rt)"

Given "A=2P, t=6"


"2P=P(1+r(6))"

"6r=1"

"r=\dfrac{1}{6}"

"r=\dfrac{100}{6}\%\approx16.6667\%"

Annual interest rate is "r=\dfrac{100}{6}\%\approx16.6667\%."